4.8 Now assume that $2 Million of Financial Services cost are related to billing and managerial reporting and $1 million are related to payroll and personnel management activities.

                a. Devise and implement a direct method cost allocation scheme that recognizes that Financial Services has two widely different functions.

                b. What, if any, additional information would be useful in completing part a?

                c. What are the cost and benefits to St. Benedict’s of creating two cost pools for Financial Services?
Answer #1

Answer :

(a) Appropriate allocation Rates

General administration (on the basis of salary dollars) Facilities (on the basis of house keeping labor house) Financial services (on the basis of patient service revenue)
Routine care 12/30 = 40% 15/22 = 68.18% 3/5 = 60%
Intensive care 5/30 = 16.67 3/22 =13.64 0.4/5 = 8%
Diagnostic services 6/30 = 20% 1.5/22 = 6.82% 0.6/5 = 12%
Other Services 7/30 = 23.33% 2.5/22 =11.36% 1/5 =20%

(b). Overhead analysis sheet

Items Basis of charges Patient service department Patient support department
Routine care Intensive care Diagnostic care Other services General administration Facilities Financial Services
Direct cost $2,000,000 $5,000,000 $3,000,000
General administration Salary dollar $800,000 $333,333 $400,000 $466,667 ($2,000,000)
Facilities House keeping labor hours $3,409,091 $681,818 $340,909 $568,182 ($5,000,000)
Financial services Service Revenue $1,800,000 $240,000 $360,000 $600,000 ($3,000,000)
Total overheads $6,009,091 $1,255,151 $1,100,909 $1,634,849 $0 $0 $0

(c) Comparison of dollar allocation

Routine care Intensive care Diagnostic services Other services
Dollar allocation in 6.3 $6,333,333 $733,333 $1,100,000 $1,833,334
Dollar allocation in 6.4 $6,009,091 $1,255,151 $1,100,909 $1,634,849
Difference $324,242 ($521,818) ($909) $198,485